A ship of mass 3.0×107kg is initially traveling at a constant speed of 12ms-1 , as it approaches land , it's engines are turned off , it decelerates uniformly and comes to rest in 30 minutes.


1
Expert's answer
2021-03-22T10:06:40-0400

Find the resistive force that makes the ship slow down. First, we need to obtain the expression for acceleration:


"a=\frac{\Delta v}{\Delta t}."

Find the force according to Newton's second law:


"F=ma=m\frac{\Delta v}{\Delta t}=3.0·10^7·\frac{12-0}{30}=12·10^6\text{ N}."